OpenStudy (anonymous):

derivative tan^-1(x/2)

OpenStudy (mandre):

You know how to use the chain rule?

OpenStudy (anonymous):

Yes

OpenStudy (mandre):

Do you know the derivative for tan^-1(x)?

OpenStudy (anonymous):

1/1+x^2

OpenStudy (mandre):

Just put x/2 into x's place and use the chain rule. Can you do it?

OpenStudy (anonymous):

Not sure how to apply chain rule to (x/2)^2

OpenStudy (mandre):

\[\frac{ 1 }{ 1+(\frac{ x }{ 2 })^{2} }.\frac{ d }{ dx }(\frac{ 1 }{ 2 }x)\]

OpenStudy (anonymous):

Alternatively, you can use drawings and implicit differentiation to determine the derivative, but knowing the formulas and rules is quicker.